OpenStudy (anonymous):

The sum of two numbers is four. The difference between ten and twice the smaller number is twelve less than four times the larger. Smaller number is ____________ Larger number is ______________

OpenStudy (anonymous):

You need to formulate and solve equations from information. Call the numbers S and L.

OpenStudy (anonymous):

First sentence says they add up to 4.

OpenStudy (anonymous):

So, you get S + L = 4 as the first step.

OpenStudy (anonymous):

Second sentence says difference between 10 and twice (2 times) the smaller number => 10 - 2S Is 12 less than four times larger number => 4L -12. So, you have 10 - 2S = 4L -12. You also have S + L =4. Substitute and solve.

OpenStudy (anonymous):

I still don't get it

OpenStudy (anonymous):

Let me ask you what part is giving you a hard time so I can explain and help better.